OpenStudy (anonymous):

A geometric sequence is defined recursively by an=4a(small n-1). the first term of the sequence is 0.5. What is the explicit formula for the nth term of the sequence?

hartnn (hartnn):

to get explicit formula you will need 2 terms, a1 = first term and r = common ratio know how to get any of these ?

OpenStudy (anonymous):

r is just u can take the 2nd term over the first term..

hartnn (hartnn):

oops, a1 = 0.5 is given :P know how to get r ?

OpenStudy (anonymous):

but idk how to find ratio out of a formula & yea a1 is given

hartnn (hartnn):

yes, or you can say any next term to current term like \(\huge r=a_n/a_{n-1}\)

hartnn (hartnn):

so, here r= .... ?

OpenStudy (anonymous):

so is the answer: a_n= 0.5*4^n-1???

hartnn (hartnn):

did you get how ? do you want to know how ?

OpenStudy (anonymous):

well the explicit 4mula for a geometric sequence is: a_n= a1*r^(n-1)..

hartnn (hartnn):

yes, a1 = 00.5 just need to find r, thats what i was explaining...

hartnn (hartnn):

that how is r=4

OpenStudy (anonymous):

wait but how can u find the ratio using that 4mula?

hartnn (hartnn):

\(\huge r=a_n/a_{n-1}\) and an = 4 a(n-1) is given

OpenStudy (anonymous):

ok so my ans is correct right?

hartnn (hartnn):

yup.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

yea?

OpenStudy (anonymous):

no don't take it like that... just leave it

OpenStudy (anonymous):

its fine rishab... just leave it ok

OpenStudy (rishabh.mission):

ok then God bye :)

OpenStudy (anonymous):

bye

hartnn (hartnn):

now that rishabh has deleted his comments, this conversation sounds weird to me :P

OpenStudy (anonymous):

he deleted his comments?

OpenStudy (anonymous):

oh wait, yeah i guess so too...

OpenStudy (rishabh.mission):

don't guess its real.... lol

OpenStudy (anonymous):

no really?

OpenStudy (rishabh.mission):

yes :)